OpenStudy (anonymous):

Help finding the exact value of this with no calculators? Step-by-step? Will give medal & fan. tan(17pi/12)

OpenStudy (anonymous):

well pi =180 degree if you didnt know

OpenStudy (anonymous):

so put that value of pi in there you get tan(255) which can be written as tan(180+75) =tan75=tan(30+45) = tan30+tan45/1-tan30tan45 sow put the values of tan30 and tan45 and find the answer

OpenStudy (anonymous):

I don't know how to get tan from anything without using a calculator...

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

memorise this table pls it is important

OpenStudy (anonymous):

Oh okay, thank you

OpenStudy (anonymous):

@ganeshie8 sorry if you are busy,i am kinda sleepy,can u pls explain this final step to her ^^

OpenStudy (anonymous):

though there is hardly anything to explain

OpenStudy (anonymous):

still @TammisaurusRex you should memorise that table

OpenStudy (anonymous):

and those formula like tan(A+B) = tanA+tanB/1-tanAtanB

OpenStudy (anonymous):

good !

OpenStudy (anonymous):

I'm writing that table down now. What do I do after I get tan(30) = 1/sqrt3 and tan(45) = 1?

OpenStudy (anonymous):

Do I just add them?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Would it be tan(30 + 45) = tan(30) + tan(45) / 1 - tan(30)tan(45) ?

OpenStudy (anonymous):

So (1/sqrt3) + 1 / 1 - (1/sqrt3)(1) (1/sqrt3) + 1 / 1 - (1/sqrt3) ?

OpenStudy (anonymous):

I just am not really sure how to add or subtract with 1/sqrt3 without a calculator...

ganeshie8 (ganeshie8):

\(\large \tan(30+45) = \dfrac{\tan 30 + \tan 45}{1- \tan 30 ~\tan 45} = \dfrac{\frac{1}{\sqrt{3}} + 1}{1 - \frac{1}{\sqrt{3}}}\)

ganeshie8 (ganeshie8):

multiply top and bottom by sqrt(3)

ganeshie8 (ganeshie8):

\(\large \dfrac{\frac{1}{\sqrt{3}} + 1}{1 - \frac{1}{\sqrt{3}}} = \dfrac{1+\sqrt{3}}{\sqrt{3} - 1}\)

ganeshie8 (ganeshie8):

rationalize the denominator

ganeshie8 (ganeshie8):

\(\large \dfrac{1+\sqrt{3}}{\sqrt{3} - 1} \times \dfrac{\sqrt{3} + 1}{\sqrt{3} + 1} = \dfrac{(\sqrt{3}+1)^2}{3-1} = \dfrac{3+2\sqrt{3} + 1}{2} = 2+\sqrt{3}\)

OpenStudy (anonymous):

Is that the exact answer?

ganeshie8 (ganeshie8):

yep ! http://www.wolframalpha.com/input/?i=tan%2817pi%2F12%29

OpenStudy (anonymous):

thank you very much!

ganeshie8 (ganeshie8):

yw